4. et X be a continuous random variable with support (0, 2) and PDF defined by f(x) = ( cx3 0 < x < 2 0 otherwise. a) Compute E[X]. b) Compute V ar(X)
Answer #1

g for et en fronde = 1 => } crºche=1 965*ck = 1 > 6. (a) => (210) = 1 => 5x16= 1 -(44=17)C=46 Lifesy of otherwise x34, OCX<2